OpenStudy (anonymous):

factor:432b^3+686q^3

OpenStudy (anonymous):

okay, you can simplify your expression first , you'll have: \[=2(216b^3 + 343q^3) = 2((6b+7q)(36b^2-42bq+49q^2))\] that's it :) I hope.

OpenStudy (anonymous):

did you get it abarnzy? :^_^